2. The body with mass m = 300 g is hung on the hook of a force gauge. Calculate the force of gravity, the force of tension exerted by the hook on the body and the weight of the body 


1
Expert's answer
2021-02-19T10:31:41-0500

The force of gravity (or the weight of the body) can be calculated as follows:


"W=mg=0.3\ kg\cdot9.8\ \dfrac{m}{s^2}=2.94\ N."

The force of tension can be found from the Newton's Second Law of Motion:


"T-mg=0,""T=mg=2.94\ N."